OpenStudy (emilyjones284):

Ill give a medal (: solve log 4x + log 13 =0

OpenStudy (anonymous):

Calculator

OpenStudy (emilyjones284):

yeah i dont understand how to do that

OpenStudy (emilyjones284):

what do i type into the calculator

OpenStudy (souvik):

log 4x=-log13 4x=10^(-log13) x=1/4*10^(-log13)